# Harmonic Tutte polynomials of matroids III

### v1

@@ 1. Introduction
In the decade prior to these results, Tutte [25] presented a two-variable polynomial of a graph, since known as the Tutte polynomial, that not only can reveal the internal structure of a graph but also can be used to determine several other graph polynomials; see also [26]. Later, Crapo [13] gave an appropriate generalization of the Tutte polynomial by defining it for matroids. Greene [17] presented a rare and unexpected relation between the weight enumerator of a linear code over finite field and the Tutte polynomial of a matroid. Moreover, using this relation, Greene [17] also gave a combinatorial proof of the MacWilliams identity for the weight enumerator of a code. Wei [27] introduced the notion of subcode weights in the study of coding theory, and Kløve [19] gave a remarkable generalization of the MacWilliams identity for these subcode weights. Many authors have studied Greene’s Theorem between the linear codes and matroids; for instance, see [10, 12]. Of note, Britz [5] generalized Greene’s Theorem with respect to the rank generating function of a matroid and the higher support weight enumerator of a linear code over a finite field. By applying this research, Britz and Shiromoto [8] generalized the Assmus-Mattson Theorem with respect to subcode supports of linear codes over finite fields; see also [4, 9, 11, 24].
@@ 2. Preliminaries
In this section, we present basic definitions and notation for linear codes and matroids that are frequently needed in this paper. We mostly follow the definitions and notations of [14, 21, 22]. We also present definitions and properties of the (discrete) harmonic functions, following [2, 3, 15] .
@@ 2.1. Discrete harmonic functions
If an element $g\in\mathbb{R}2^{E}$ is equal to $\widetilde{f}$ for some $f\in\mathbb{R}E_{d}$, then we say that $g$ has degree $d$. The differentiation $\gamma$ is the operator on $\mathbb{R}E_{d}$ defined by linearity from the identity
@@
Then
@@
If $X=\{1,3,4\}$, then $\widetilde{f}(X)=a_{2}+a_{3}+a_{6}$. Suppose that $f\in\operatorname{Harm}_{2}(4)$; then $\gamma(f)=0$, so
$a_{1}+a_{2}+a_{3}=a_{1}+a_{4}+a_{5}=a_{2}+a_{4}+a_{6}=a_{3}+a_{5}+a_{6}=0.$
Solving the above equations, we get
$\operatorname{Harm}_{2}(4)\ni f=a_{1}\{1,2\}+a_{2}\{1,3\}+(-a_{1}-a_{2})\{1,4\}\\ +(-a_{1}-a_{2})\{2,3\}+a_{2}\{2,4\}+a_{1}\{3,4\}.$
@@ 2.2. Linear codes
is called the $r$-th higher weight enumerator of $C$ and satisfies the following MacWilliams-type identity stated in [18]:
@@
The following two remarks provide important and useful identities.
Remark 2.3.
@@ 3. Coding theory associated to harmonic functions
In this section, we introduce the harmonic generalization of higher weight enumerators for codes over finite fields.
@@ 3.1. Harmonic higher weight enumerators
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$ where $d\neq 0$.
@@
If $\deg f=0$, then the harmonic $r$-th higher weight enumerator $W_{C,f}^{(r)}(x,y)$ is the $r$-th higher weight enumerator $W_{C}^{(r)}(x,y)$.
Example 3.1.
Let $C$ be the binary linear code of length $3$ with codewords
$(0,0,0),(0,0,1),(1,1,0),(1,1,1)\,.$
We consider $f\in\operatorname{Harm}_{1}(3)$ follows:
$\operatorname{Harm}_{1}(3)\ni f=a\{1\}+b\{2\}-(a+b)\{3\}.$
Then by direct computation, we find the harmonic weight enumerator of $C$ associated to $f$ to be
$W_{C,f}^{(1)}=-(a+b)x^{2}y+(a+b)xy^{2}=(xy)^{1}Z_{C,f}^{(1)},$
where $Z_{C,f}^{(1)}=(a+b)(y-x)$.
Theorem 3.1 (MacWilliams type identity).
Let $W_{C,f}^{(r)}(x,y)$ be the harmonic $r$-th weight enumerator of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ associated to $f\in\operatorname{Harm}_{d}(n)$. Then
$W_{C,f}^{(r)}(x,y)=(xy)^{d}Z_{C,f}^{(r)}(x,y),$
where $Z_{C,f}^{(r)}$ is a homogeneous polynomial of degree $n-2d$, and satisfies
$Z_{C,f}^{(\ell)}\left({x+(q^{j}-1)y},{x-y}\right).$
In order to give a proof of the above theorem, we first present some lemmas and theorems on harmonic functions and their associated polynomials $Z_{C,f}^{(r)}$.
@@
We have from the above equality that if $|J|>n-d$, then $\widetilde{f}(J)=0$.
From the above discussion and the first part of Theorem 3.1, we see that
$Z_{C,f}^{(r)}(x,y)=\sum_{i=d}^{n-d}A_{i,f}^{(r)}x^{n-i-d}y^{i-d}.$
@@
Then we have the following relation between $B_{t,f}^{(r)}(C)$ and $A_{i,f}^{(r)}(C)$ as follows:
@@
It is immediate from (2) and Remark 3.4 that $B_{t,f}^{(r)}(C)=0$ for $0\leq t<d$ and $n-d<t\leq n$. We now focus on $t$ with $d\leq t\leq n-d$. By Remark 3.4 and Remark 2.3,
@@
$=(-1)^{d}\sum_{i=d}^{n-t}\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\sum_{j=0}^{d}\binom{n-i-d+j}{n-i-t-d+j}\sum_{\begin{subarray}{c}Z\in E_{d},\\ |\operatorname{Supp}(D)\cap Z|=j\end{subarray}}f(Z)$
By Lemma 3.1 and the identity [16, 3.49]
$\sum_{j=0}^{d}(-1)^{j}\binom{x-j}{t}\binom{d}{j}=\binom{x-d}{t-d}\,,$
it follows that
@@ 3.2. Harmonic extended weight enumerators
Let $m$ be a positive integer. Then any $[n,k]$ code $C$ over $\mathbb{F}_{q}$ can be extended to a linear code $C{[q^{m}]}:=C\otimes\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q^{m}}$ by considering all the $\mathbb{F}_{q^{m}}$-linear combinations of the codewords in $C$. We call $C{[q^{m}]}$ the $\mathbb{F}_{q^{m}}$-extension code of $C$. Clearly, $\dim_{\mathbb{F}_{q}}C(X)=\dim_{\mathbb{F}_{q^{m}}}C[q^{m}](X)$ for all $X\subseteq E$ (see [18]).
@@
Now we have the following MacWilliams type identity for the $\mathbb{F}_{q^{m}}$-extended code over $\mathbb{F}_{q^{m}}$. The proof of this identity is similar to the proofs of [2, Theorem 2.1] and [24, Theorem 2.1]. One can also follow an alternative way given in [12] to prove this identity. We therefore omit the proof of the following theorem.
@@
Let $W_{C,f}(x,y;q^{m})$ be the harmonic $q^{m}$-extended weight enumerator of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ associated to $f\in\operatorname{Harm}_{d}(n)$. Then
$W_{C,f}(x,y;q^{m})=(xy)^{d}Z_{C,f}(x,y;q^{m}),$
where $Z_{C,f}(x,y;q^{m})$ is a homogeneous polynomial of degree $n-2d$ that satisfies
@@
$\sum\limits_{X\in E_{t}}\widetilde{f}(X)(q^{m})^{\ell(X)}$ $=\sum\limits_{X\in E_{t}}\widetilde{f}(X)(((q^{m})^{\ell(X)}-1)+1)$
$=\sum\limits_{X\in E_{t}}\widetilde{f}(X)(B_{X}^{q^{m}}(C)+1)$
@@
The proof of the above proposition is similar to the proof of [12, Proposition 3.8] and is therefore omitted. Using the above proposition, we get the following relation.
@@
Proof.
By using Proposition 3.2, we can prove the theorem with an argument similar to that used to prove [12, Proposition 3.9]. ∎
@@ 3.3. The link between the two weight enumerators
Let $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ be the $\mathbb{F}_{q}$-linear space of $m\times n$ matrices with entries in $\mathbb{F}_{q}$. Assume that $M\in\operatorname{Mat}(m\times n,\mathbb{F}_{q})$. Then for an $\mathbb{F}_{q}$-linear code $C$ of length $n$, $S_{m,n}(C)$ denotes the linear subspace of $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ such that rows of each matrix are in $C$. Then by [18, Proposition 5.27], there is an isomorphism from $C{[q^{m}]}$ into $S_{m,n}(C)$. For a detailed discussion, we refer the reader to [18, 23].
@@
Let $\mathbf{u}\in C{[q^{m}]}$ and $M\in S_{m,n}(C)$ be the corresponding $m\times n$ matrix under a given isomorphism $C[q^{m}]\to S_{m,n}(C)$. Let $D_{M}\subseteq C$ be the subcode generated by the rows of $M$. Then $\operatorname{supp}(\mathbf{u})=\operatorname{Supp}(D_{M})$, and hence $\operatorname{wt}(\mathbf{u})=\operatorname{wt}(D_{M})$ (see [18]). Therefore,
@@
By Theorem 3.2, Remarks 2.4 and 2.3 and Proposition 3.3,
@@ 4. A Greene-type theorem
$T(M,f;x,y):=\sum_{X\subset E}\widetilde{f}(X)(x-1)^{\rho(E)-\rho(X)}(y-1)^{|X|-\rho(X)}.$
It is not so difficult to show the relation between the Tutte polynomial of a matroid and its dual associated to a harmonic function:
@@
The proof follows very similar argument to those in [12, Theorem 4.2]. We therefore leave the details to the reader. ∎
@@
Lemma 3.1 and Remark 3.4 implies that $Z_{C,f}^{(r)}(x,y)$ is a polynomial. By Theorems 3.6 and 3.3, we can write
@@
We consider $f\in\operatorname{Harm}_{1}(5)$ follows:
$\operatorname{Harm}_{1}(5)\ni f=a\{1\}+b\{2\}+c\{3\}+d\{4\}-(a+b+c+d)\{5\}.$
By direct calculation, we have

### v2

# Harmonic higher and extended weight enumerators

@@ 1. Introduction
In the decade prior to these results, Tutte [25] presented a two-variable polynomial of a graph, since known as the Tutte polynomial, that not only can reveal the internal structure of a graph but also can be used to determine several other graph polynomials; see also [26]. Later, Crapo [12] gave an appropriate generalization of the Tutte polynomial by defining it for matroids. Greene [16] presented a rare and unexpected relation between the weight enumerator of a linear code over finite field and the Tutte polynomial of a matroid. Moreover, using this relation, Greene [16] also gave a combinatorial proof of the MacWilliams identity for the weight enumerator of a code. Wei [27] introduced the notion of subcode weights in the study of coding theory, and Kløve [18] gave a remarkable generalization of the MacWilliams identity for these subcode weights. Many authors have studied Greene’s Theorem between the linear codes and matroids; for instance, see [11]. Of note, Britz [5] generalized Greene’s Theorem with respect to the rank generating function of a matroid and the higher support weight enumerator of a linear code over a finite field. By applying this research, Britz and Shiromoto [8] generalized the Assmus-Mattson Theorem with respect to subcode supports of linear codes over finite fields; see also [4, 9, 10, 24].
@@ 2. Preliminaries
In this section, we present basic definitions and notation for linear codes and matroids that are frequently needed in this paper. We mostly follow the definitions and notation of [13, 21, 22]. We also present definitions and properties of (discrete) harmonic functions, following [2, 3, 14] .
@@ 2.1. Discrete harmonic functions
Note that $\widetilde{f}(\emptyset)=f(\emptyset)$ when $d=0$, and that $\widetilde{f}(\emptyset)=0$ otherwise. If an element $g\in\mathbb{R}2^{E}$ is equal to $\widetilde{f}$ for some $f\in\mathbb{R}E_{d}$, then we say that $g$ has degree $d$. The differentiation operator $\gamma$ on $\mathbb{R}E_{d}$ is defined by linearity from the identity
@@
If $X=\{1,3,4\}$, then $\widetilde{f}(X)=a_{2}+a_{3}+a_{6}$. Now applying differential operator $\gamma$ on $f$, we have
@@
It follows that $f\in\operatorname{Harm}_{2}(4)$ if and only if $a_{3}=a_{4}=-a_{1}-a_{2}$, $a_{5}=a_{2}$ and $a_{6}=a_{1}$.
@@ 2.2. Linear codes
is called the $r$-th higher weight enumerator of $C$ and satisfies the following MacWilliams-type identity stated in [17, Theorem 5.14]:
@@
The following two lemmas provide important and useful identities. The second lemma follows immediately from [19, Theorem 25.2].
Lemma 2.1.
@@ 3. Coding theory associated to harmonic functions
Bachoc [2] first defined the notion of harmonic weight enumerators for binary codes and presented a MacWilliams-type identity for these enumerators. Later, several articles extended this concept for linear codes over finite fields [3, 24] as well as over finite rings [7]. In this section, we introduce the harmonic generalization of higher (resp. extended) weight enumerators for codes over finite fields.
Definition 3.1.
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$. Let $f\in\operatorname{Harm}_{d}(n)$, where $d\neq 0$. Then the harmonic weight enumerator of $C$ associated to $f$ is defined as follows:
$W_{C,f}(x,y):=\sum_{i=0}^{n}A_{i,f}(C)x^{n-i}y^{i},$
where
$A_{i,f}(C):=\sum_{\mathbf{u}\in C,\operatorname{wt}(\mathbf{u})=i}\widetilde{f}(\operatorname{supp}(\mathbf{u})).$
@@ 3.1. Harmonic higher weight enumerators
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$, where $d\neq 0$.
@@
If $\deg f=0$, then $W_{C,f}^{(r)}(x,y)=f(\emptyset)\,W_{C}^{(r)}(x,y)$.
@@
Moreover, we have from the above equality that if $|J|>n-d$, then $\widetilde{f}(J)=0$.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Then it is immediate from Remark 3.4 that the harmonic $r$-th higher weight enumerator $W_{C,f}^{(r)}(x,y)$ of $C$ associated to $f\in\operatorname{Harm}_{d}(n)$ for any $r$, $1\leq r\leq k$ can be written as:
${W_{C,f}^{(r)}(x,y)=(xy)^{d}Z_{C,f}^{(r)}(x,y),}$
where $Z_{C,f}^{(r)}$ is a homogeneous polynomial of degree $n-2d$ of the form as follows:
${Z_{C,f}^{(r)}(x,y)=\sum_{i=d}^{n-d}A_{i,f}^{(r)}(C)x^{n-i-d}y^{i-d}.}$
Example 3.1.
Let $C$ be the binary linear code of length $3$ with codewords
$(0,0,0),(0,0,1),(1,1,0),(1,1,1)\,.$
We consider $f\in\operatorname{Harm}_{1}(3)$, where $f=a\{1\}+b\{2\}-(a+b)\{3\}$. Then by direct computation, we find the harmonic weight enumerator of $C$ associated to $f$ to be
$W_{C,f}^{(1)}=-(a+b)x^{2}y+(a+b)xy^{2}=(xy)^{1}Z_{C,f}^{(1)},$
where $Z_{C,f}^{(1)}=(a+b)(y-x)$.
Theorem 3.1 (MacWilliams type identity).
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C,f}^{(\ell)}\left({x+(q^{j}-1)y},{x-y}\right).$
To pave the way for the proof of Theorem 3.1 that we give in Section 4, we begin by presenting some propositions and theorems related to $Z_{C,f}^{(r)}(x,y)$.
@@
Then we have the following relation between $B_{t,f}^{(r)}(C)$ and $A_{i,f}^{(r)}(C)$:
@@
It is immediate from (2) and Remark 3.4 that $B_{t,f}^{(r)}(C)=0$ for $0\leq t<d$ and $n-d<t\leq n$. We now focus on $t$ with $d\leq t\leq n-d$. By Remark 3.4, we have
@@
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{j=0}^{d}\binom{n-i-d+j}{n-i-t-d+j}\sum_{\begin{subarray}{c}Z\in E_{d},\\ |\operatorname{Supp}(D)\cap Z|=j\end{subarray}}f(Z)\,.$
Using inclusion-exclusion to count the $t$-element subsets of an $x$-element set that include some fixed $d$ elements gives the identity
(5) $\sum_{j=0}^{d}(-1)^{j}\binom{x-j}{t}\binom{d}{j}=\binom{x-d}{t-d}\,.$
By Lemma 3.1 and (5), it follows that
@@ 3.2. Harmonic extended weight enumerators
Let $m$ be a positive integer. Then any $[n,k]$ code $C$ over $\mathbb{F}_{q}$ can be extended to a linear code $C{[q^{m}]}:=C\otimes\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q^{m}}$ by considering all the $\mathbb{F}_{q^{m}}$-linear combinations of the codewords in $C$. We call $C{[q^{m}]}$ the $\mathbb{F}_{q^{m}}$-extension code of $C$. Clearly, $\dim_{\mathbb{F}_{q}}C(X)=\dim_{\mathbb{F}_{q^{m}}}C[q^{m}](X)$ for all $X\subseteq E$ (see [17, p. 34]).
@@
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Then by Remark 3.4 one can easily observe that, for any positive integer $m$, the harmonic $q^{m}$-extended weight enumerator $W_{C,f}(x,y;q^{m})$ of $C$ associated to $f\in\operatorname{Harm}_{d}(n)$ can be written as
${W_{C,f}(x,y;q^{m})=(xy)^{d}Z_{C,f}(x,y;q^{m}),}$
where $Z_{C,f}(x,y;q^{m})$ is the following homogeneous polynomial of degree $n-2d$:
${Z_{C,f}(x,y;q^{m})=\sum_{i=d}^{n-d}A_{i,f}^{q^{m}}(C)x^{n-i-d}y^{i-d}.}$
Now we have the following harmonic analogue of the MacWilliams identity for $\mathbb{F}_{q^{m}}$-extended code. Bachoc [2] and Tanabe [24] presented the harmonic generalizations of the MacWilliams identity for binary and non-binary codes, respectively. Later, an alternative approach of the proof of this identity was given in [11].
@@
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
@@
Following the approach given in [11], we provide the proof of Theorem 3.3 in Section 4. In order to do that we need to present some necessary propositions and theorems related to $Z_{C,f}(x,y;q^{m})$.
@@
$\sum\limits_{X\in E_{t}}\widetilde{f}(X)(q^{m})^{\ell(X)}$ $={\sum\limits_{X\in E_{t}}\widetilde{f}(X)(B_{X}^{q^{m}}(C)+1)}$
@@
The proof of the above proposition is similar to the proof of [11, Proposition 3.8] and is therefore omitted. Using the above proposition, we have the extended analogue of [11, Proposition 3.9] as follows. Since the proof is straightforward, we leave it to the reader.
@@ 3.3. The link between the two weight enumerators
Let $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ be the $\mathbb{F}_{q}$-linear space of $m\times n$ matrices with entries in $\mathbb{F}_{q}$. Then for an $\mathbb{F}_{q}$-linear code $C$ of length $n$, $S_{m,n}(C)$ denotes the linear subspace of $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ such that rows of each matrix are in $C$. Then by [17, Proposition 5.27], there is an isomorphism from $C{[q^{m}]}$ into $S_{m,n}(C)$. For a detailed discussion, we refer the reader to [17, 23].
@@
Let $\mathbf{u}\in C{[q^{m}]}$ and $M\in S_{m,n}(C)$ be the corresponding $m\times n$ matrix under a given isomorphism $C[q^{m}]\to S_{m,n}(C)$. Let $D_{M}\subseteq C$ be the subcode generated by the rows of $M$. Then $\operatorname{supp}(\mathbf{u})=\operatorname{Supp}(D_{M})$, and hence $\operatorname{wt}(\mathbf{u})=\operatorname{wt}(D_{M})$ (see [17, Lemma 5.4]). Therefore,
@@
By Theorem 3.2, Lemmas 2.2 and 2.1 and Proposition 3.3,
@@ 4. A Greene-type theorem
$T(M,f;x,y):={\sum_{X\subseteq E}}\widetilde{f}(X)(x-1)^{\rho(E)-\rho(X)}(y-1)^{|X|-\rho(X)}.$
By Remark 3.4, the following relation between the harmonic Tutte polynomial of a matroid and that of its dual is immediately clear:
@@
The proof follows very similar argument to those in [11, Theorem 4.2]. We therefore leave the details to the reader. ∎
@@
Let $C$ be an $[n,k]$ code, and $M_{C}$ be its matroid. Let $k^{\perp}$ be the dimension of $C^{\perp}$. Then by Theorem 4.1,
$(-1)^{d}$ $q^{-(k-d)m}Z_{C,f}({x+(q^{m}-1)y},{x-y};q^{m})$
$=\frac{(-1)^{2d}}{q^{(k-d)m}}(q^{m}y)^{k-d}(x-y)^{n-k-d}T\left(M_{C},f;\dfrac{x}{y},\dfrac{x+(q^{m}-1)y}{x-y}\right)$
$=(x-y)^{n-k-d}y^{k-d}T\left(M_{C},f;\dfrac{x}{y},\dfrac{x+(q^{m}-1)y}{x-y}\right)$
$=(-1)^{d}(x-y)^{(n-k)-d}y^{n-(n-k)-d}T\left(M_{C^{\perp}},f;\dfrac{x+(q^{m}-1)y}{x-y},\dfrac{x}{y}\right)$
$=(-1)^{d}(x-y)^{k^{\perp}-d}y^{n-k^{\perp}-d}T\left(M_{C^{\perp}},f;\dfrac{x+(q^{m}-1)y}{x-y},\dfrac{x}{y}\right)$
$=Z_{C^{\perp},f}(x,y)\,.\qed$
Proof of Theorem 3.1.
By Theorems 3.6 and 3.3, we can write
@@
We consider $f\in\operatorname{Harm}_{1}(5)$, where $f=a\{1\}+b\{2\}+c\{3\}+d\{4\}-(a+b+c+d)\{5\}$. By direct calculation, we have
@@
The following remarks concerning harmonic generalization of the MacWilliams identity for $r$-th higher weight enumerators stated in Theorem 3.1 might be helpful:
Remark 4.1.
Using the binomial series expansion for $i=d,\ldots,n-d$, we obtain the identity
$(x+(q^{j}-1)y)^{n-d-i}(x-y)^{i-d}$ $=\sum_{p=d}^{n-d}K_{p}(i-d;n-2d,q^{j})x^{n-d-p}y^{p-d},$
where $K_{p}(i-d;n-2d,q^{j})$ is known as the Krawtchouk polynomial. For any positive integer $n>2d$, it is defined as follows (see [23]):
Remark 4.2.
Since $A_{i,f}^{(0)}(C^{\perp})=0$, we can write $A_{i,f}^{(r)}(C^{\perp})$ explicitly for $i=d,\ldots,n-d$ and $r\neq 0$ as
$A_{i,f}^{(r)}(C^{\perp})=(-1)^{d}\sum_{\ell=1}^{r}\,$ $\sum_{j=\ell}^{r}(-1)^{r-j}\frac{q^{{r-j\choose 2}-j(r-j)-\ell(j-\ell)-j(k-d)}}{[r-j]_{q}[j-\ell]_{q}}$
$\sum_{p=d}^{n-d}K_{p}(i-d;n-2d,q^{j})A_{p,f}^{(\ell)}(C).$